OpenStudy (luigi0210):

A farmer wishes to fence in a rectangular field. If north-south fencing costs $3 a yard and east-west fencing costs $2 a yard, what are the dimensions of the field of maximum area that can be fenced in for $600?

pooja195 (pooja195):

Let the north-south length be W and the east-west length be L. Target function: A = WxL, where A is the area of the fence. Restriction: 3(2W) + 2(2L) = 600 Subsitution: A = W(150 - 1.5W) Differentiate A with respect to W, A' = 150 - 1.5W - 1.5W Solve A' = 0 for W, W = 50 yard L = 75 yard

pooja195 (pooja195):

do u get this

OpenStudy (luigi0210):

Yes, thank you :)